# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v2

@@ 12.6.1 Breaking New Ground: The First-Ever 4-Way Liver Exchange
After several entries into the liver exchange pool, the number of patient-donor pairs reached 12.132132132Although I use the phrase “patient-donor pairs,” it is not uncommon for patients to have multiple willing donors. Therefore, the intended meaning is patients with one or more donors. Upon receiving a new set of data, I couldn’t believe my eyes. As always, I reviewed it before Ünver ran the software he’d coded for the system. Immediately, I spotted a 4-way exchange alongside a separate 2-way exchange. Remarkably, it was already possible to match at least half of the patients in the pool! By then, even a 3-way liver exchange was unprecedented.133133133As it turns out, with the support of Alex Chan—a graduate student at Stanford Business School, a world-first 3-way liver exchange was carried out in Pakistan a few months earlier on March 17, 2022. We became aware of this development in December 2022 when its report became available online in JAMA Surgery (Salman et al., 2023).
@@ 12.6.3 The Pilot Phase in the First Year
12.6.4 Milestones Post-July 2023: World’s First 5-Way, 6-Way, and 7-Way Exchanges
@@
The public announcement marked the conclusion of the pilot phase of the BBS-LPE System. Over the following sixteen months, it facilitated an additional 177 LDLTs—averaging approximately 11 transplants per month—and bringing the total to 192 by the first week of December 2024. This rate surpassed the outcomes of even the most active liver exchange systems by an order of magnitude. Indeed, prior to our system, only one center, Medanta in India, had managed to perform as many as 10 liver exchange transplants in a single calendar year. This period was marked by one groundbreaking achievement after another for the system.
The Malatya team performed the world’s first 5-way, 6-way, and 7-way exchanges in October 2023, January 2024, and July 2024, respectively (see Figure 9 for a depiction of the world’s first 7-way liver exchange). They became the first team to surpass 100 liver exchange transplants in March 2024—less than two years after the system’s launch in June 2022—and the first team to perform 100 liver exchange transplants in a calendar year in October 2024.
As of December 2024, the BBS-LPE System accounts for both 7-way exchanges, all four 6-way, all four 5-way, and all eight 4-way conducted globally, along with sixteen of the nineteen 3-way liver exchanges ever performed, firmly establishing itself as the global leader in liver exchange.
@@ 12.6.5 Comparison with Other Prominent Liver Exchange Programs
The BBS-LPE program, though not the first of its kind, has rapidly emerged as the global leader in liver exchange within just two years of its deployment. To understand the key factors driving this success, let us first compare its outcomes with those of other prominent liver exchange programs worldwide.
In 2023, despite a catastrophic earthquake that disrupted life across the region for months, the BBS-LPE System facilitated 64 LDLTs, accounting for 27.7% of the 231 LDLTs performed by the Malatya team that year. Building on this achievement, the system enabled 116 LDLTs in the first eleven months of 2024—an unprecedented 47.3% of the 245 LDLTs performed during that period.
Cumulatively, the system’s contributions include 2 liver exchange transplants in 2021, 10 in 2022, 64 in 2023, and 116 so far in 2024, bringing the total to 192. To put this into perspective, prior to the BBS-LPE System’s deployment, the global total of reported liver exchange transplants over two decades since 2002 was fewer than 250. In just two years, the Malatya team has performed nearly 80% of the liver exchange transplants reported globally over that time.
@@
In the U.S., liver exchange programs are currently in their early phases. As of December 2024, a total of 172 liver exchange transplants have been reported in the U.S., an all-time high of 50 of them occurring in 2023 (OPTN, 2024a). Most activity is concentrated at two centers. The liver exchange program at the University of Texas Health San Antonio reported 23 liver exchange transplants from 2019 through October 2023 (Klair et al., 2024). These transplants were conducted through four 2-way exchanges, six 2-way non-directed donor chains, and one 3-way non-directed donor chain. Similarly, the program at the University of Pittsburgh Medical Center reported 20 liver exchange transplants from January 2019 through July 2021 (Gunabushanam et al., 2022), achieved through three 2-way exchanges and seven 2-way non-directed donor chains.
@@ 12.7 Malatya’s Journey to Global Leadership: Implementing Best Practices
Our two decades of experience in kidney exchange have revealed three key factors that boost the efficacy of living-donor organ exchange programs. The first is the ability to incorporate altruistic non-directed donors (NDDs) into the system and multiply their impact through NDD chains. This approach is how the National Kidney Registry became the leading kidney exchange platform in the U.S. (see Section 7.9.1), and how Canada developed one of the most successful national kidney exchange systems in the world (see Section 7.9.2). Unfortunately, because altruistic donation is illegal in Turkey, this was not an option for the Malatya team.
@@ 13.1 Leveling the Playing Field in School Choice
To build a compelling case against this mechanism in favor of the Deferred Acceptance (DA) or Top Trading Cycles (TTC) mechanisms, Yan Chen and I conducted a laboratory experiment involving all three (Chen and Sönmez, 2006).134134134Building on Chen and Sönmez (2006), a substantial body of literature has employed lab experiments to study preference manipulation in school choice mechanisms. For a comprehensive review of this literature and other behavioral studies in education market design, see Rees-Jones and Shorrer (2023). For a broader review of behavioral studies in market design, see Chen et al. (2021). In our experimental setting, both DA and TTC outperformed the Boston mechanism.
@@
Fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. The set of Nash equilibrium outcomes of this game is equal to the set of outcomes that satisfy individual rationality, non-wastefulness, and no justified envy.135135135Equivalently, by Proposition 4, the set of Nash equilibrium outcomes of this game is equal to the set of stable outcomes of the induced many-to-one matching market.
@@
Partitioning the students into strategic and sincere groups, fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. There exists a Nash equilibrium outcome in this game that is weakly preferred by each strategic student to the outcome of the DA mechanism in truthful dominant strategies.136136136With the BPS data used in Pathak and Sönmez (2008), the Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism is virtually unique under all partitions of students into strategic and sincere.
@@ 13.2 Opting for Obscurity: The Curious Case of Seattle School Choice
In Pathak and Sönmez (2013), we identified several indicators that the BW amendment was not well understood, even though a strategy-proof mechanism should have enabled the school district to advise families that submitting truthful preferences was a safe strategy. Confusion about the BW amendment even surfaced during a 2001 deposition of the school board president in a court challenge to Seattle’s school choice plan.137137137Page 58 of the U.S. Court of Appeals for the Ninth Circuit, No. 01–35450, Parents Involved in Community Schools vs. Seattle School District No.1, 2001.
@@
If reducing transportation costs was indeed a factor in this reform, it would exemplify an objective that does not meet my definition of “legitimate.” Rather than directly addressing school choice, were decision-makers attempting to limit the mobility it introduced indirectly by reverting to the Boston mechanism? This rationale would also be consistent with the decision to keep the strategy-proofness of the school choice mechanism opaque for nearly a decade—from 1999 to 2008.
@@ 13.3 Nudging School Choices: Taiwan’s Deduction Mechanism
Despite these relatively modest deductions, it is evident that—partly due to the mechanism’s explicit penalties for lower-ranked choices—TD mechanisms other than DA are all susceptible to preference manipulation. Thus, of all TD mechanisms, only DA satisfies the axiom of strategy-proofness. Similarly, none of the TD mechanisms, except for DA, satisfy the axiom of no justified envy.138138138It’s important to note that the school priority rankings induced by the effective scores under the TD mechanism are artifacts of the mechanism, not inherent to the problem. Therefore, no justified envy is defined based on the original student scores from the CAE.
@@
In this context, the TD mechanism’s deductions act as a nudge, pushing families toward more conservative school rankings.139139139While nudges are typically viewed as tools to guide individuals toward beneficial decisions—i.e., paternalistic nudges—in this case, they represent a form of nudging “for bad,” as discussed in Glaeser (2006); Thaler (2015); Schmidt and Engelen (2020).
@@ 13.4 Framing Institutions: How Descriptions Shape Their Adoption
Taiwan’s experience with its school choice mechanism illustrates how the framing of an institution can significantly influence both participant behavior and its acceptance by central planners, revealing the potential role of behavioral economics while exposing the limitations of traditional mechanism design in some practical contexts. In this abstract framework, a mechanism consists of a message space for each agent and an outcome function that maps message profiles to outcomes. It assumes that neither the process by which this mapping is derived nor how it is communicated affects participant behavior or the mechanism’s appeal to central planners. However, Taiwan’s case suggests that these assumptions may not always hold in real-world settings.140140140This observation aligns with the recent work of Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) on mechanism descriptions that make strategy-proofness easier to observe. Given a direct mechanism $\phi$, an individual $i$, and the reported types (preferences, in our context) of all other individuals, the menu for $i$ under $\phi$ is constructed by applying the mechanism across all possible types for $i$. Hammond (1979) shows that a direct mechanism is strategy-proof if and only if each agent always receives one of their best outcomes from their menu. Focusing on strategy-proof mechanisms, Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) suggest that the strategy-proofness of a mechanism can be made evident to participants through its menu descriptions, with experimental evidence indicating that truthful preference revelation increases under such descriptions. Their work highlights that, irrespective of the mechanism’s specific formulation, how it is communicated can significantly influence participant behavior.
@@
Consider the first formulation. As illustrated in Example 7, women are disadvantaged in the second batch of positions since their highest-merit-score members have already received assignments, while the highest-merit-score men remain. However, this disadvantage may not be obvious to the layperson.141141141Indeed, indicative of a conservative estimate of this oversight, an online experiment by Pathak et al. (2023) reveals that about 40% of subjects believe the order in which positions are allocated is irrelevant. In contrast, in the second formulation, it is clear that the benefits for women under this affirmative action policy are conditional, only activating if women do not receive 30 positions without affirmative action. Thus, the relative modesty of this policy is more apparent in the second formulation. Therefore, a decision maker seeking to adopt a more potent policy might be more inclined to choose the minimum guarantee choice rule when presented with its original formulation, perhaps overlooking its modesty, than if they were presented with the adjustment-based formulation.
@@ 14.1 Conservatism within Minimalist Market Design
Compared to proponents of economics imperialism, minimalist market design offers a far less intrusive approach when applying economic formalism and methodologies to non-economic fields and practical applications.142142142For a similar perspective against economic imperialism, see also Rodrik (2021), who emphasizes that while economists’ approaches can complement those of other scholarly disciplines, they cannot substitute the often qualitative methods employed in these fields. It adopts an interdisciplinary stance without insisting on neoclassical principles.143143143Importantly, minimalist market design does not reject the main principles of neoclassical economics; it simply does not assume them. By adhering to its first two pillars, minimalist market design largely avoids the criticisms raised by Santos and other critics of economics imperialism, which may explain its success in interdisciplinary collaborations. While often operating outside the boundaries of traditional economics—extending into areas such as education, constitutional law, and medical sciences—it does not promote profit-seeking or strategic behavior. Instead, as explored throughout this monograph, the minimalist approach seeks to eliminate the need for such behavior.
Some critics, drawing on the ”performativity thesis” of Callon (1998), argue that design economists create a performative framework in real-world implementations (Mirowski and Nik-Khah, 2007; Santos, 2011; Breslau, 2013; Hitzig, 2018, 2020). They contend that this process shapes participants and institutions to align with theoretical models, effectively making reality conform to economists’ constructs.144144144Two additional concerns related to the performativity thesis are the normative gap and algorithmic bias. In applications where social justice issues are important, Hitzig (2020) highlights a concern that market designers’ formulation of social goals and methods can diverge from policymakers’ objectives, leading to what she calls the normative gap. This occurs when the goals implemented by economic engineers do not fully align with the intentions of policymakers. Meanwhile, algorithmic bias arises as markets increasingly rely on rules and algorithms to achieve objectives, with some researchers warning that these designs might unfairly advantage certain groups (Hitzig, 2018; van Basshuysen, 2022). When executed properly, minimalist market design can address both issues. In cases where there is a unique mechanism satisfying stakeholders’ goals, neither a normative gap nor algorithmic bias arises. When multiple solutions are possible, best practices involve maintaining informed neutrality among reasonable normative positions (Li, 2017) and providing policymakers with a representative list of minimalist interventions. This approach ensures that the design aligns with the institution’s mission and stakeholders’ objectives, thereby minimizing both the normative gap and algorithmic bias.

### v3

@@ 12.6.1 Breaking New Ground: The First-Ever 4-Way Liver Exchange
After several entries into the liver exchange pool, the number of patient-donor pairs reached 12.136136136Although I use the phrase “patient-donor pairs,” it is not uncommon for patients to have multiple willing donors. Therefore, the intended meaning is patients with one or more donors. Upon receiving a new set of data, I couldn’t believe my eyes. As always, I reviewed it before Ünver ran the software he’d coded for the system. Immediately, I spotted a 4-way exchange alongside a separate 2-way exchange. Remarkably, it was already possible to match at least half of the patients in the pool! By then, even a 3-way liver exchange was unprecedented.137137137As it turns out, with the support of Alex Chan—a graduate student at Stanford Business School, a world-first 3-way liver exchange was carried out in Pakistan a few months earlier on March 17, 2022. We became aware of this development in December 2022 when its report became available online in JAMA Surgery (Salman et al., 2023).
@@ 12.6.3 The Pilot Phase in the First Year
12.6.4 Milestones Post-Pilot Phase: World’s First 5-Way, 6-Way, and 7-Way Exchanges
@@
The public announcement marked the end of the BBS-LPE System’s pilot phase. Over the next seventeen months, it facilitated an additional 183 LDLTs—with an average of more than ten transplants per month—and brought the total to 198 by the end of 2024. This pace surpassed the outcomes of even the most active liver exchange programs by an order of magnitude. Before the BBS-LPE System’s introduction, only one center—Medanta in India—had performed as many as ten liver exchange transplants in a single calendar year. In contrast, this period witnessed a series of unprecedented milestones:
  • World’s First Multi-Way Exchanges: The Malatya team performed the world’s first 5-way, 6-way, and 7-way exchanges in October 2023, January 2024, and July 2024, respectively (see Figure 9 for a depiction of the first 7-way liver exchange).
  • Record-Breaking Overall Volume: By March 2024—less than two years after the system’s June 2022 launch—the Malatya team became the first in the world to surpass 100 cumulative liver exchange transplants.
  • Unmatched Annual Performance: In October 2024, the team became the first to exceed 100 liver exchange transplants within a single calendar year.
As of December 2024, the BBS-LPE System accounts for both 7-way exchanges, all four 6-way, all four 5-way, and all eight 4-way conducted globally, along with eighteen of the twenty-one 3-way liver exchanges ever reported, firmly establishing itself as the global leader in liver exchange.
@@ 12.6.5 Comparison with Other Prominent Liver Exchange Programs
The BBS-LPE program, though not the first of its kind, has rapidly emerged as the global leader in liver exchange in just two years since its deployment. To understand the key factors behind this success, it is helpful to compare its outcomes with those of other prominent liver exchange programs worldwide.
In 2023, despite a catastrophic earthquake that disrupted life across the region for months, the BBS-LPE System facilitated 64 LDLTs, accounting for 27.7% of the 231 LDLTs performed by the Malatya team that year. Building on this achievement, the system enabled 122 LDLTs in 2024—an unprecedented 45.2% of the 270 LDLTs performed during this period.
Cumulatively, the system facilitated 2 liver exchange transplants in 2021, 10 in 2022, 64 in 2023, and 122 in 2024, bringing the total to 198 by the end of 2024. To put this into perspective, the global total of reported liver exchange transplants between 2002 and 2022 was fewer than 250 prior to the BBS-LPE System’s deployment. Remarkably, in just two years, the Malatya team facilitated approximately 80% of the global total of reported liver exchange transplants accumulated over the two decades preceding its launch.
@@
In the U.S., liver exchange programs are currently in their early phases. As of December 2024, a total of 173 liver exchange transplants have been reported in the U.S., an all-time high of 50 of them occurring in 2023 (OPTN, 2024a). Most activity is concentrated at two centers. The liver exchange program at the University of Texas Health San Antonio reported 23 liver exchange transplants from 2019 through October 2023 (Klair et al., 2024). These transplants were conducted through four 2-way exchanges, six 2-way non-directed donor chains, and one 3-way non-directed donor chain. Similarly, the program at the University of Pittsburgh Medical Center reported 20 liver exchange transplants from January 2019 through July 2021 (Gunabushanam et al., 2022), achieved through three 2-way exchanges and seven 2-way non-directed donor chains.
@@ 12.7 Malatya’s Journey to Global Leadership: Implementing Best Practices
Our two decades of experience in kidney exchange have revealed three key factors that boost the efficacy of living-donor exchange programs. The first is the ability to incorporate altruistic non-directed donors (NDDs) into the system and amplify their impact through NDD chains. This approach enabled the National Kidney Registry to become the leading kidney exchange platform in the U.S. (see Section 7.9.1) and helped the U.K. and Canada establish two of the most successful national kidney exchange systems in the world (see Section 7.9.2). Unfortunately, because altruistic donation is illegal in Turkey, this was not an option for the Malatya team.
@@ 13.1 Leveling the Playing Field in School Choice
To build a compelling case against this mechanism in favor of the Deferred Acceptance (DA) or Top Trading Cycles (TTC) mechanisms, Yan Chen and I conducted a laboratory experiment involving all three (Chen and Sönmez, 2006).138138138Building on Chen and Sönmez (2006), a substantial body of literature has employed lab experiments to study preference manipulation in school choice mechanisms. For a comprehensive review of this literature and other behavioral studies in education market design, see Rees-Jones and Shorrer (2023). For a broader review of behavioral studies in market design, see Chen et al. (2021). In our experimental setting, both DA and TTC outperformed the Boston mechanism.
@@
Fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. The set of Nash equilibrium outcomes of this game is equal to the set of outcomes that satisfy individual rationality, non-wastefulness, and no justified envy.139139139Equivalently, by Proposition 4, the set of Nash equilibrium outcomes of this game is equal to the set of stable outcomes of the induced many-to-one matching market.
@@
Partitioning the students into strategic and sincere groups, fix a school choice problem and consider the preference revelation game induced by the Boston mechanism. There exists a Nash equilibrium outcome in this game that is weakly preferred by each strategic student to the outcome of the DA mechanism in truthful dominant strategies.140140140With the BPS data used in Pathak and Sönmez (2008), the Nash equilibrium outcome of the preference revelation game induced by the Boston mechanism is virtually unique under all partitions of students into strategic and sincere.
@@ 13.2 Opting for Obscurity: The Curious Case of Seattle School Choice
In Pathak and Sönmez (2013), we identified several indicators that the BW amendment was not well understood, even though a strategy-proof mechanism should have enabled the school district to advise families that submitting truthful preferences was a safe strategy. Confusion about the BW amendment even surfaced during a 2001 deposition of the school board president in a court challenge to Seattle’s school choice plan.141141141Page 58 of the U.S. Court of Appeals for the Ninth Circuit, No. 01–35450, Parents Involved in Community Schools vs. Seattle School District No.1, 2001.
@@
If reducing transportation costs was indeed a factor in this reform, it would exemplify an objective that does not meet my definition of “legitimate.” Rather than directly addressing school choice, were the decision makers attempting to limit the mobility it introduced indirectly by reverting to the Boston mechanism? This rationale would also be consistent with the decision to keep the strategy-proofness of the school choice mechanism opaque for nearly a decade—from 1999 to 2008.
@@ 13.3 Nudging School Choices: Taiwan’s Deduction Mechanism
Despite these relatively modest deductions, it is evident that—partly due to the mechanism’s explicit penalties for lower-ranked choices—TD mechanisms other than DA are all susceptible to preference manipulation. Thus, of all TD mechanisms, only DA satisfies the axiom of strategy-proofness. Similarly, none of the TD mechanisms, except for DA, satisfy the axiom of no justified envy.142142142It’s important to note that the school priority rankings induced by the effective scores under the TD mechanism are artifacts of the mechanism, not inherent to the problem. Therefore, no justified envy is defined based on the original student scores from the CAE.
@@
The deductions in the TD mechanism, in general, do not function purely as a nudge, since they directly affect student options and incentives. However, suppose a central planner were to describe the Boston mechanism using its TD mechanism formulation with deductions. If students respond to this formulation with more conservative strategies than they do with the standard formulation, then the deductions, in this case, would function purely as a nudge.
In this sense, beyond their direct effect on choices and incentives, the TD mechanism’s deductions also act as a nudge, encouraging families to adopt more conservative school rankings.143143143While nudges are typically viewed as tools to guide individuals toward beneficial decisions—i.e., paternalistic nudges—in this case, they represent a form of nudging “for bad,” as discussed in Glaeser (2006); Thaler (2015); Schmidt and Engelen (2020).
@@ 13.4 Framing Institutions: How Descriptions Shape Their Adoption
Taiwan’s experience with its school choice mechanism illustrates how the framing of an institution can significantly influence both participant behavior and its acceptance by central planners, revealing the potential role of behavioral economics while exposing the limitations of traditional mechanism design in some practical contexts. In this abstract framework, a mechanism consists of a message space for each agent and an outcome function that maps message profiles to outcomes. It assumes that neither the process by which this mapping is derived nor how it is communicated affects participant behavior or the mechanism’s appeal to central planners. However, Taiwan’s case suggests that these assumptions may not always hold in real-world settings.144144144This observation aligns with the recent work of Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) on mechanism descriptions that make strategy-proofness easier to observe. Given a direct mechanism $\phi$, an individual $i$, and the reported types (preferences, in our context) of all other individuals, the menu for $i$ under $\phi$ is constructed by applying the mechanism across all possible types for $i$. Hammond (1979) shows that a direct mechanism is strategy-proof if and only if each agent always receives one of their best outcomes from their menu. Focusing on strategy-proof mechanisms, Gonczarowski et al. (2023, 2024) and Katuscak and Kittsteiner (2024) suggest that the strategy-proofness of a mechanism can be made evident to participants through its menu descriptions, with experimental evidence indicating that truthful preference revelation increases under such descriptions. Their work highlights that, irrespective of the mechanism’s specific formulation, how it is communicated can significantly influence participant behavior.
@@
Consider the first formulation. As illustrated in Example 7, women are disadvantaged in the second batch of positions since their highest-merit-score members have already received assignments, while the highest-merit-score men remain. However, this disadvantage may not be obvious to the layperson.145145145Indeed, indicative of a conservative estimate of this oversight, an online experiment by Pathak et al. (2023) reveals that about 40% of subjects believe the order in which positions are allocated is irrelevant. In contrast, in the second formulation, it is clear that the benefits for women under this affirmative action policy are conditional, only activating if women do not receive 30 positions without affirmative action. Thus, the relative modesty of this policy is more apparent in the second formulation. Therefore, a decision maker seeking to adopt a more potent policy might be more inclined to choose the minimum guarantee choice rule when presented with its original formulation, perhaps overlooking its modesty, than if they were presented with the adjustment-based formulation.
@@ 14.1 Conservatism within Minimalist Market Design
Compared to proponents of economics imperialism, minimalist market design offers a far less intrusive approach when applying economic formalism and methodologies to non-economic fields and practical applications.146146146For a similar perspective against economic imperialism, see also Rodrik (2021), who emphasizes that while economists’ approaches can complement those of other scholarly disciplines, they cannot substitute the often qualitative methods employed in these fields. It adopts an interdisciplinary stance without insisting on neoclassical principles.147147147Importantly, minimalist market design does not reject the main principles of neoclassical economics; it simply does not assume them. By adhering to its first two pillars, minimalist market design largely avoids the criticisms raised by Santos and other critics of economics imperialism, which may explain its success in interdisciplinary collaborations. While often operating outside the boundaries of traditional economics—extending into areas such as education, constitutional law, and medical sciences—it does not promote profit-seeking or strategic behavior. Instead, as explored throughout this monograph, the minimalist approach seeks to eliminate the need for such behavior.
Some critics, drawing on the ”performativity thesis” of Callon (1998), argue that design economists create a performative framework in real-world implementations (Mirowski and Nik-Khah, 2007; Santos, 2011; Breslau, 2013; Hitzig, 2018, 2020). They contend that this process shapes participants and institutions to align with theoretical models, effectively making reality conform to economists’ constructs.148148148Two additional concerns related to the performativity thesis are the normative gap and algorithmic bias. In applications where social justice issues are important, Hitzig (2020) highlights a concern that market designers’ formulation of social goals and methods can diverge from policymakers’ objectives, leading to what she calls the normative gap. This occurs when the goals implemented by economic engineers do not fully align with the intentions of policymakers. Meanwhile, algorithmic bias arises as markets increasingly rely on rules and algorithms to achieve objectives, with some researchers warning that these designs might unfairly advantage certain groups (Hitzig, 2018; van Basshuysen, 2022). When executed properly, minimalist market design can address both issues. In cases where there is a unique mechanism satisfying stakeholders’ goals, neither a normative gap nor algorithmic bias arises. When multiple solutions are possible, best practices involve maintaining informed neutrality among reasonable normative positions (Li, 2017) and providing policymakers with a representative list of minimalist interventions. This approach ensures that the design aligns with the institution’s mission and stakeholders’ objectives, thereby minimizing both the normative gap and algorithmic bias.